# PF3plat: Pose-Free Feed-Forward 3D Gaussian Splatting

### v1

@@ 3.2.3 Camera Pose Refinement
In this module, we further refine camera poses to enhance reconstruction and view synthesis quality. Initially, we replace the estimated relative poses $\mathcal{P}_{ij}$ with newly computed camera poses $\hat{\mathcal{P}}_{ij}$ derived from following the similar process in coarse alignment and using the previously obtained refined depths $\hat{\mathcal{D}}_{i}$. We then introduce a learnable camera pose refinement module that estimates rotation and translation offsets. To streamline this process, we first utilize a fully differentiable transformation synchronization operation that takes $\hat{\mathcal{P}}_{ij}$ and $\mathcal{C}_{ij}$ as inputs. Using power iterations (El Banani et al., 2023), this operation efficiently recovers the absolute poses $\hat{\mathcal{P}}_{i}$ prior to the refinement module.
@@
where $\Delta R,\Delta t$ are the pose offsets, and $E_{\mathrm{pos}})$ is positional embedding.
@@
Cost Volume Construction and Aggregation.
@@
Specifically, given the current pose estimates $\hat{\mathcal{P}}_{i}$, we build a multi-view stereo cost volume $\mathcal{C}_{i}^{\mathrm{multi}}\in\mathbb{R}^{h\times w\times K}$ following the plane-sweeping approach (Yao et al., 2018; Chen et al., 2021). For each of the $K$ depth candidates within specified near and far ranges along the epipolar lines, we compute matching scores using cosine similarity (Cho et al., 2022; 2021; Hong et al., 2024). Subsequently, to guide the depth localization along the epipolar lines, we construct a guidance cost volume $\mathcal{C}_{i}^{\mathrm{guide}}$ (Li et al., 2023), where each spatial location is represented by a one-hot vector indicating the depth candidate closest to the monocular depth estimate. The constructed $\mathcal{C}_{i}^{\mathrm{multi}}$ and $\mathcal{C}_{i}^{\mathrm{guide}}$ undergo cross-attention to update the multi-view cost volume $\hat{\mathcal{C}}_{i}^{\mathrm{multi}}$.
@@
From the updated multi-view cost volume $\hat{\mathcal{C}}_{i}^{\mathrm{multi}}$, we apply a softmax function (Xu et al., 2022; Hong et al., 2023) along the $K$ dimension to obtain a matching distribution. We then extract the maximum value from this distribution to derive a confidence score, $S_{\mathrm{geo}}$, which assesses the quality of the estimated camera pose and depth. Formally, this is defined as:
$S_{i}^{\mathrm{geo}}=\max_{k\in{\{1,2,...,K\}}}\mathrm{softmax}(\hat{\mathcal{C}}_{i}^{\mathrm{multi}})(k).$ (3)
These confidence scores assess the reliability of the predicted 3D Gaussian centers, where high confidence indicates accurate localization and low confidence suggests potential inaccuracies due to noise or misalignment. To condition the prediction of Gaussian parameters, opacity, covariance, and color, we incorporate $S_{\mathrm{geo}}$ as additional input.
@@
Finally, using the inputs $[I_{i},\hat{\mathcal{D}}_{i},F_{i},S_{i}^{\mathrm{geo}}]$, we compute the opacity $\sigma_{i}$ through small convolutional layers, derive the covariances from the estimated rotations and scales, and obtain the color from the estimated spherical harmonic (SH) coefficients. A key idea of our approach is that $S^{\mathrm{geo}}$ enables supervision signals to flow from the Gaussian parameters back to the depth and pose estimates. This feedback loop enhances the accuracy of both depth and pose estimations, resulting in more consistent and reliable 3D reconstructions.
@@ 3.3 Loss Function
By making both depth and camera pose refinement modules learnable and differentiable, our network leverages supervision signals from the pixel-aligned 3D Gaussians and the novel view synthesis task (Zhou et al., 2017) to refine these estimates during training. Specifically, we combine photometric loss, defined as the L2 loss between the rendered and target images, as well as SSIM (Wang et al., 2004) loss $\mathcal{L}_{\text{SSIM}}$ and LPIPS (Zhang et al., 2018) loss $\mathcal{L}_{\text{LPIPS}}$ to form our reconstruction loss $\mathcal{L}_{\text{img}}$.
@@
We identify that provided good coarse alignments, RGB loss is sufficient, but with larger baselines, the training process starts to destabilize. Moreover, one remaining issue with learning solely from the photometric loss is that the gradients are mainly derived from pixel intensity differences, which suffer in textureless regions. To remedy these, we enforce that corresponding points in the set of images $\{I\}_{i=1}^{N}$ lie on the same object surface, drawing from principles of multi-view geometry Hartley & Zisserman (2003).
Formally, using the estimated depths $\hat{D}$, the camera poses $\hat{\mathcal{P}}$, and the correspondence sets $\mathcal{M}$, we can define a geometric consistency loss that penalizes deviations from the multi-view geometric constraints. For each correspondence $(p,q)\in\mathcal{M}_{ij}$ between images $I_{i}$ and $I_{j}$, we compute the 3D point from the pixel $p$ and its estimated depth $\hat{D}_{i}(p)$ using the camera intrinsics. We then transform this to the coordinate frame of $I_{j}$ using the relative pose $\hat{\mathcal{P}}_{ij}$ and project it back onto the image plane to obtain the predicted correspondence $\tilde{p}$. This is defined as $\mathcal{L}_{\mathrm{2D-3D}}=\sum_{(p,q)\in\mathcal{M}}\varphi(\tilde{p},q),$ where $\varphi(\cdot)$ denotes huber loss. By integrating this loss into the learning pipeline, we improve the robustness of the model in regions with low texture or significant viewpoint changes, ensuring that the estimated pose and depth are consistent with the underlying 3D structure of the scene.
@@
Combining the three loss functions, we define our final objective function: $\mathcal{L}_{\mathrm{img}}+\mathcal{L}_{\mathrm{2D-3D}}+\lambda_{\mathrm{3D-3D}}\mathcal{L}_{\mathrm{3D-3D}},$ where we set $\lambda_{\mathrm{3D-3D}}=0.05$
@@ 4.1 Implementation Details
In this work, we assume intrinsic parameters are given, as they are generally available from modern devices (Arnold et al., 2022). We compute attentions using Flash Attention (Dao et al., 2022), and for the Gaussian rasterizer, we follow the method described in (Charatan et al., 2023). Our model is trained on a single NVIDIA A6000 GPU for 40,000 iterations using the Adam optimizer (Kingma, 2014), with a learning rate set to $2\times 10^{-4}$ and a batch size of 5, which takes approximately two days. For training on the RealEstate10K and ACID datasets, we gradually increase the number of frames between $I_{1}$ and $I_{2}$ as training progresses, initially setting the frame distance to 45 and gradually increasing it to 75. For the DL3DV dataset, we start with a frame distance of 5 and increase it to 10. The target view is randomly sampled within this range. We train under the assumption of $N=2$ and render $\hat{I}_{t}$. Our code is implemented using PyTorch (Paszke et al., 2017). Additional implementation details can be found in the supplementary material. The code and pretrained weights will be made publicly available.
@@
Ours 23.732 0.251 0.702 0.008 23.719 0.250 0.702 0.009 23.935 0.246 0.708 0.008 23.647 0.253 0.695 0.008
4.2 Experimental Setting
Datasets.
We train and evaluate our method on three large-scale datasets: RealEstate10K (Zhou et al., 2018), a collection of both indoor and outdoor scenes; ACID (Liu et al., 2021), a dataset focusing on outdoor coastal scenes; and DL3DV (Ling et al., 2024), which includes diverse real-world indoor and outdoor environments. For RealEstate10K, due to some unavailable videos on YouTube, we use a subset of the full dataset, comprising a training set of 21,618 scenes and a test set of 7,200 scenes. For ACID, we train on 10,935 scenes and evaluate on 1,893 scenes. Lastly, for DL3DV, which features longer video sequences across 10,510 different scenes, we use 2,000 scenes for training and a standard benchmark set of 140 scenes for testing (Ling et al., 2024).
Baselines.
Following Hong et al. (2024), we evaluate our method on two tasks: novel-view synthesis and camera pose estimation. For novel view synthesis, we compare our approach against established generalized NeRF and 3DGS variants, including PixelNeRF (Yu et al., 2021), Du et al. (2023), PixelSplat (Charatan et al., 2023), and MVSPlat (Chen et al., 2024). It is important to note that these methods assume ground-truth (GT) camera poses during inference, so we include them only for reference. Our primary comparisons focus on existing pose-free generalized novel view synthesis methods, such as DBARF (Chen & Lee, 2023), FlowCAM (Smith et al., 2023), and CoPoNeRF (Hong et al., 2024).
For camera pose estimation, we first evaluate against correspondence-based pose estimation methods (e.g., COLMAP), including SP+SG (DeTone et al., 2018; Sarlin et al., 2020), PDC-Net+(Truong et al., 2023a), DUSt3R(Wang et al., 2023), and MASt3R (Leroy et al., 2024). Additionally, we compare with direct pose regression methods such as 8ViT (Rockwell et al., 2022), RelPose (Zhang et al., 2022), and MicKey (Barroso-Laguna et al., 2024). Our main comparisons, however, are with existing pose-free approaches, including DBARF (Chen & Lee, 2023), FlowCAM (Smith et al., 2023), and CoPoNeRF (Hong et al., 2024).
@@ 4.2 Experimental Setting
  Task Method Avg Small Medium Large
@@
Pose Estimation 8ViT 12.59 6.881 90.12 88.65 12.60 6.860 91.46 91.50 12.17 6.552 82.48 82.92 12.77 7.214 91.85 88.92
@@
  Task Method Avg Small Medium Large
@@
Pose Estimation 8ViT 4.568 1.312 88.433 88.961 8.466 3.151 88.421 88.958 4.325 1.564 90.555 90.799 2.280 0.699 86.580 87.559
@@
CoPoNeRF 3.283 1.134 22.809 14.502 3.548 1.129 23.689 11.289 2.573 1.169 21.401 10.656 3.455 1.129 22.935 10.588
Ours 4.125 1.776 27.727 13.903 4.011 1.604 27.786 13.840 3.667 1.604 26.343 13.622 4.412 1.891 28.330 14.092
Table 2: Pose Estimation Performance on RealEstate-10K and ACID. Gray entries indicate methods not trained on the same dataset due to the absence of ground truth data (e.g., depth and correspondence), making them incomparable.
@@
For the evaluation on RealEstate-10K and ACID, we follow the protocol outlined by Hong et al. (2024), where evaluation is conducted using unposed triplet images $(I_{1},I_{2},I_{t})$. The test set is divided into three groups, small, middle, and large, based on the extent of overlap between $I_{1}$ and $I_{2}$. This allows the model’s performance to be assessed under varying levels of difficulty, reflecting different real-world scenarios. For the relatively new DL3DV dataset, we introduce a new evaluation protocol. For each scene, we select two context images, $I_{1}$ and $I_{2}$, by skipping frames at intervals of 5 and 10, creating two groups per scene, each representing small and large overlap cases. We then randomly select three target images from the sequence between the context images, similar to the above protocol.
For evaluation metrics, we use standard image quality measures, PSNR, SSIM, LPIPS, and MSE, for novel view synthesis. For camera pose estimation, we compute the geodesic rotation error and angular difference in translation, as commonly done in classical methods (Nistér, 2004; Melekhov et al., 2017). These errors are measured by comparing the ground truth relative pose $P^{\mathrm{GT}}_{12}$ and our estimated pose $\hat{\mathcal{P}}_{12}$. Our statistical analysis reports both the average and median errors, with the median providing robustness against outliers.
@@
Ours 19.355 0.280 0.611 4.698 2.475 10.692 6.869 22.105 0.211 0.706 3.353 1.604 9.407 6.334
Table 3: Novel View Synthesis and Pose Estimation Performance on DL3DV. We include MVSPlat for reference only.
@@ 4.3 Experimental Results
Tab.1 summarizes the performance for the novel view synthesis task, while Tab.2 reports the results for pose estimation. From the results in Tab. 1, our method significantly outperforms previous pose-free generalizable methods (Chen & Lee, 2023; Smith et al., 2023; Hong et al., 2024), setting a new state-of-the-art across these benchmarks. Furthermore, compared to the previous state-of-the-art, CoPoNeRF, our method achieves a 2.8 dB improvement in PSNR, demonstrating superior reconstruction quality and robustness. Additionally, our approach demonstrates superior pose estimation performance on RealEstate-10K; however, we observe that Hong et al. (2024) achieves lower pose errors on the ACID dataset. This discrepancy may be attributed to the larger scale of scenes, such as coastal landscapes and sky views, which complicates our refinement process and poses challenges for our depth network in estimating the metric depth of the scene. Nonetheless, this limitation is mitigated by our method’s superior novel view image quality and the fact that Hong et al. (2024) utilizes ground-truth poses for supervision, providing robust guidance for large-scale environments. Moreover, the ACID dataset includes numerous dynamic scenes, which are beyond the scope of our current focus. Consequently, our method may be less effective in estimating poses for dynamic scenes compared to other approaches.
@@
While RealEstate-10K and ACID encompass a variety of indoor and outdoor scenes, RealEstate-10K predominantly includes indoor environments, whereas ACID features numerous dynamic scenes. To comprehensively evaluate our method across a broader spectrum of real-world scenarios, we further assess it on the recently released DL3DV dataset (Ling et al., 2024). The results are summarized in Table 3. From these results, we observe that our method outperforms CoPoNeRF (Hong et al., 2024) by over 4 dB in large-overlap scenarios and by 3.8 dB in small-overlap scenarios, highlighting the superior accuracy and robustness of our approach in handling diverse and complex environments. This highlights the effectiveness of our method in managing varied scene and object types, reinforcing its applicability for practical novel view synthesis tasks.

### v2

@@ 3.2.3 Camera Pose Refinement
Here, we further refine camera poses to enhance reconstruction and view synthesis quality. Initially, we replace the estimated relative poses $\mathcal{P}_{ij}$ with newly computed camera poses $\hat{\mathcal{P}}_{ij}$ derived from following the similar process in coarse alignment and using the previously obtained refined depths $\hat{\mathcal{D}}_{i}$. We then introduce a learnable camera pose refinement module that estimates rotation and translation offsets. To streamline this process, we first utilize a fully differentiable transformation synchronization operation that takes $\hat{\mathcal{P}}_{ij}$ and $\mathcal{C}_{ij}$ as inputs. Using power iterations (El Banani et al., 2023), this operation efficiently recovers the absolute poses $\hat{\mathcal{P}}_{i}$ prior to the refinement module.
@@
where $\Delta R,\Delta t$ are the pose offsets, and $E_{\mathrm{pos}}$ is positional embedding.
@@
Multi-View and Guidance Cost Volume Construction and Aggregation.
@@
Specifically, given the current pose estimates $\hat{\mathcal{P}}_{i}$, we build a multi-view stereo cost volume $\mathcal{C}_{i}^{\mathrm{multi}}\in\mathbb{R}^{h\times w\times K}$ following the plane-sweeping approach (Yao et al., 2018; Chen et al., 2021, 2024; An et al., 2024). For each of the $K$ depth candidates within specified near and far ranges along the epipolar lines, we compute matching scores using cosine similarity (Cho et al., 2022, 2021; Hong et al., 2024). Subsequently, to guide the depth localization along the epipolar lines, we construct a guidance cost volume $\mathcal{C}_{i}^{\mathrm{guide}}$ (Li et al., 2023), where each spatial location is represented by a one-hot vector indicating the depth candidate closest to the monocular depth estimate. We finally obtain an aggregated cost volume ${\mathcal{C}}_{i}^{\mathrm{agg}}$ by feeding $\mathcal{C}_{i}^{\mathrm{multi}}$ and $\mathcal{C}_{i}^{\mathrm{guide}}$ to a series of cross-attention layers to update the multi-view cost volume guided by a guidance cost volume. We define the process as following:
${\mathcal{C}}_{i}^{\mathrm{agg}}=\mathcal{T}_{agg}(\mathcal{C}_{i}^{\mathrm{multi}},\mathcal{C}_{i}^{\mathrm{guide}}),$ (3)
where $\mathcal{T}(\cdot)$ is a deep transformer architecture that computes cross-attention between inputs.
@@
Using the aggregated cost volume $\mathcal{C}_{i}^{\mathrm{agg}}$, we apply a softmax function (Xu et al., 2022; Hong et al., 2023) along the $K$ dimension to obtain a matching distribution. We then extract the maximum value from this distribution to derive a confidence score, $S_{\mathrm{geo}}$, which assesses the quality of the estimated camera pose and depth. Formally, this is defined as:
$S_{i}^{\mathrm{geo}}=\max_{k\in{\{1,2,...,K\}}}\mathrm{softmax}({\mathcal{C}}_{i}^{\mathrm{agg}})(k).$ (4)
These confidence scores assess the reliability of the predicted 3D Gaussian centers, where high confidence indicates accurate localization and low confidence suggests potential inaccuracies due to noise or misalignment. To condition the prediction of Gaussian parameters, such as opacity, covariance, and color, we incorporate $S_{\mathrm{geo}}$ as additional input.
@@
Finally, using the inputs $[I_{i},\hat{\mathcal{D}}_{i},F_{i},S_{i}^{\mathrm{geo}}]$, we compute the opacity $\sigma_{i}$ through small convolutional layers, derive the covariances from the estimated rotations and scales, and obtain the color from the estimated SH coefficients. A key idea of our approach is that $S^{\mathrm{geo}}$ enables supervision signals to flow from the Gaussian parameters back to the depth and pose estimates. This feedback loop enhances the accuracy of both depth and pose estimations, resulting in more consistent and reliable 3D reconstructions.
@@ 3.3 Loss Function
We combine photometric loss, defined as the L2 loss between the rendered and target images, as well as SSIM (Wang et al., 2004) loss $\mathcal{L}_{\text{SSIM}}$ and LPIPS (Zhang et al., 2018) loss $\mathcal{L}_{\text{LPIPS}}$ to form our reconstruction loss $\mathcal{L}_{\text{img}}$.
@@
We identify that provided good coarse alignments, RGB loss is sufficient, as similarly observed in (Ye et al., 2024), but with larger baselines, the training process starts to destabilize. Moreover, one remaining issue with learning solely from the photometric loss is that the gradients are mainly derived from pixel intensity differences, which suffer in textureless regions. To remedy these, we enforce that corresponding points in the set of images $\{I\}_{i=1}^{N}$ lie on the same object surface, drawing from principles of multi-view geometry (Hartley & Zisserman, 2003).
Formally, using the estimated depths $\hat{D}$, the camera poses $\hat{\mathcal{P}}$, and the correspondence sets $\mathcal{M}$, we can define a geometric consistency loss that penalizes deviations from the multi-view geometric constraints. For each correspondence $(p,q)\in\mathcal{M}_{ij}$ between images $I_{i}$ and $I_{j}$, we compute the 3D point from the pixel $p$ and its estimated depth $\hat{D}_{i}(p)$ using the camera intrinsics. We then transform this to the coordinate frame of $I_{j}$ using the relative pose $\hat{\mathcal{P}}_{ij}$ and project it back onto the image plane to obtain the predicted correspondence $\tilde{p}$. This is defined as $\mathcal{L}_{\mathrm{2D-3D}}=\sum_{(p,q)\in\mathcal{M}}\varphi(\tilde{p},q),$ where $\varphi(\cdot)$ denotes huber loss.
@@
Figure 3: Qualitative Comparison on RealEstate-10K and ACID. Given two context views (a) and (b), we compare novel view rendering results.
@@
Combining the three loss functions, we define our final objective function: $\mathcal{L}_{\mathrm{img}}+\mathcal{L}_{\mathrm{2D-3D}}+\lambda_{\mathrm{3D-3D}}\mathcal{L}_{\mathrm{3D-3D}},$ where we set $\lambda_{\mathrm{3D-3D}}=0.05.$
@@ 4.1 Implementation Details
We compute attentions using Flash Attention (Dao et al., 2022), and for the Gaussian rasterizer, we follow the method described in (Charatan et al., 2023). Our model is trained on 4 NVIDIA A100 GPU for 50,000 iterations using the Adam optimizer (Kingma, 2014), with a learning rate set to $8\times 10^{-4}$ and a batch size of 9 per each GPU, which takes approximately two days. For training on the RealEstate10K and ACID datasets, we gradually increase the frame distance between $I_{1}$ and $I_{2}$ as training progresses, initially setting the frame distance to 15 and gradually increasing it to 75. For the DL3DV dataset, we start with a frame distance of 5 and increase it to 10. The target view is randomly sampled within this range. The code and pretrained weights will be made publicly available.
4.2 Experimental Setting
Datasets.
We train and evaluate our method on three large-scale datasets: RealEstate10K (Zhou et al., 2018), a collection of both indoor and outdoor scenes; ACID (Liu et al., 2021), a dataset focusing on outdoor coastal scenes; and DL3DV (Ling et al., 2024), which includes diverse real-world indoor and outdoor environments. For RealEstate10K, due to some unavailable videos on YouTube, we use a subset of the full dataset, comprising a training set of 21,618 scenes and a test set of 7,200 scenes. For ACID, we train on 10,935 scenes and evaluate on 1,893 scenes. Lastly, for DL3DV, we train on 10,510 different scenes and evaluate on the standard benchmark set of 140 scenes for testing (Ling et al., 2024).
Baselines.
Following (Hong et al., 2024), we evaluate our method on two tasks: novel-view synthesis and camera pose estimation. For novel view synthesis, we compare our approach against established generalized NeRF and 3DGS variants, including PixelNeRF (Yu et al., 2021), (Du et al., 2023), PixelSplat (Charatan et al., 2023), and MVSPlat (Chen et al., 2024). It is important to note that these methods assume GT camera poses during training and inference, so we include them only for reference. Our primary comparisons focus on existing pose-free generalized novel view synthesis methods, such as DBARF (Chen & Lee, 2023), FlowCAM (Smith et al., 2023), and CoPoNeRF (Hong et al., 2024).
@@ 4.2 Experimental Setting
Ours 25.640 0.204 0.784 0.006 25.882 0.205 0.788 0.006 25.998 0.211 0.790 0.006 25.321 0.201 0.778 0.006
For camera pose estimation, we evaluate against correspondence-based pose estimation methods (e.g., SfM), including SP+SG (DeTone et al., 2018; Sarlin et al., 2020), PDC-Net+ (Truong et al., 2023a), DUSt3R (Wang et al., 2023), and MASt3R (Leroy et al., 2024). Due to absence of GT depth in our datasets, we leverage their pre-trained weights and direct comparison is avoided. Additionally, we compare with direct pose regression methods such as 8ViT (Rockwell et al., 2022), RelPose (Zhang et al., 2022). Our main comparisons include pose-free view synthesis approaches, including DBARF (Chen & Lee, 2023), FlowCAM (Smith et al., 2023), and CoPoNeRF (Hong et al., 2024).
Table 2: Pose Estimation Performance on RealEstate-10K and ACID. Gray entries indicate methods that were not trained on the same dataset due to missing ground-truth data (e.g., depth or correspondences), making them not directly comparable to ours. In other words, we cannot train our approach on their dataset, nor can they train theirs on ours. *: We also include a MASt3R variant that omits iterative pose optimization (roughly taking $\sim$ 10s) and instead relies on a PnP solver to maintain consistency with other baselines.
@@
  Task Additional Info. Method Avg Small Medium Large
@@
MASt3R* 3.392 1.455 24.346 8.997 5.048 1.954 14.232 5.472 2.045 1.261 19.574 8.581 1.576 1.059 42.385 28.390
Pose Estimation Pose 8ViT 12.59 6.881 90.12 88.65 12.60 6.860 91.46 91.50 12.17 6.552 82.48 82.92 12.77 7.214 91.85 88.92
@@
  Task Additional Info. Method Avg Small Medium Large
@@
MASt3R* 3.988 1.438 45.376 25.917 4.458 1.461 45.328 27.233 3.214 1.364 45.903 27.540 4.062 1.473 45.160 24.870
Pose Estimation Pose 8ViT 4.568 1.312 88.433 88.961 8.466 3.151 88.421 88.958 4.325 1.564 90.555 90.799 2.280 0.699 86.580 87.559
@@
Pose CoPoNeRF 3.283 1.134 22.809 14.502 3.548 1.129 23.689 11.289 2.573 1.169 21.401 10.656 3.455 1.129 22.935 10.588
Ours 2.691 1.113 20.319 9.366 2.551 0.998 22.888 8.889 1.989 0.788 17.884 8.887 3.112 1.338 19.887 9.887
@@
For evaluation, we follow the protocol outlined by (Hong et al., 2024) using unposed triplet images $(I_{1},I_{2},I_{t})$, with the test set divided into small, middle, and large based on the extent of overlap between $I_{1}$ and $I_{2}$. We also provide multi-view evaluation in Tab. 5(c). We provide further details in the supplementary materials.
Table 3: Novel View Synthesis and Pose Estimation Performance on DL3DV. We include PixelSplat and MVSplat for reference only.
@@
Ours 19.822 0.248 0.651 4.338 2.339 9.998 6.532 22.668 0.198 0.723 3.448 1.598 9.338 6.177
@@ 4.3 Experimental Results
Tab.1 summarizes the performance for the novel view synthesis task, while Tab.2 reports the results for pose estimation. From the results in Tab. 1, our method significantly outperforms previous pose-free generalizable methods (Chen & Lee, 2023; Smith et al., 2023; Hong et al., 2024), achieving a 4.0 dB improvement to CoPoNeRF in PSNR, demonstrating superior reconstruction quality and robustness. We also provide qualitative comparison in Fig. 3. Additionally, our approach also demonstrates superior pose estimation performance on both datasets, even surpassing (Hong et al., 2024) that trains its network with GT camera poses; however, we observe that compared to MASt3R, we slightly fall behind on ACID dataset. This discrepancy may be attributed to the larger scale of scenes, such as coastal landscapes and sky views, or dynamic scenes in ACID which complicates our refinement process and poses challenges for our depth network in estimating the metric depth of the scene.
@@
While RealEstate-10K and ACID encompass a variety of indoor and outdoor scenes, RealEstate-10K predominantly includes indoor environments, whereas ACID features numerous dynamic scenes. To comprehensively evaluate our method across a broader spectrum of real-world scenarios, we further assess it on the recently released DL3DV dataset (Ling et al., 2024). The results are summarized in Table 3. From these results, we observe that our method outperforms CoPoNeRF (Hong et al., 2024) by over 5 dB in large-overlap scenarios and by 4 dB in small-overlap scenarios, highlighting the superior accuracy and robustness of our approach in handling diverse and complex environments. This highlights the effectiveness of our method in managing varied scene and object types, reinforcing its applicability for practical view synthesis tasks.